What is the Y intersect and slope of a graph that contains (4,-6) and (9,2)
Answer by ilana(305)

First of all, it's "y-intercept".
So, we can get this information by getting the equation of the line in y-intercept form, or y=mx+b.
We have 2 points, so we can find the slope=m. m= ((2)-(-6))/((9)-(4)). I used many parenthesis here so I do not confuse my positives and negatives.
So m=(2+6)/(9-4)=8/5.
Now, we have a slope and a point, so we can write the equation in point-slope form, y-y1=m(x-x1). The point we use for (x1,y1) is insignificant, so I will use (4,-6), giving y+6=8/5(x-4), so y+6=8/5x-32/5, so y=8/5x-(32/5+6), so y=8/5x-(32/5+30/5), so y=8/5x-62/5.
We now have a line in the form y=mx+b, which we know makes m the slope and b the y-intercept, so the slope is 8/5 and the y-intercept is -62/5=-12.4